OpenStudy (anonymous):

Which of the following is a boxplot for this distribution? 54, 59, 35, 41, 46, 25, 47, 60, 54, 46, 49, 46, 41, 34, 22

OpenStudy (amistre64):

since you have provided no pictures to choose from, youll still need to detemrine the quartiles for the box

OpenStudy (amistre64):

its simplest if you sort the values from small to large finding a quartile is then a positional game

OpenStudy (amistre64):

lol, they really dont like you do they.

OpenStudy (amistre64):

still have to do it the long way them

OpenStudy (amistre64):

a quartile (a quarte, or simply 1/4) of a set of "k" elements is k/4 q0 = 0*k/4 q1 = 1*k/4 q2 = 2*k/4 q3 = 3*k/4 q4 = 4*k/4 all these qs led to a position in the lineup

OpenStudy (amistre64):

so, how many numbers do we have in the set?

OpenStudy (anonymous):

15

OpenStudy (amistre64):

so, 15/4 is a quartile, find the positional values that relate to the qs

OpenStudy (anonymous):

amistre64 help on my question please?

OpenStudy (amistre64):

the wolf does a clean sort function: http://www.wolframalpha.com/input/?i=sort+%7B54%2C+59%2C+35%2C+41%2C+46%2C+25%2C+47%2C+60%2C+54%2C+46%2C+49%2C+46%2C+41%2C+34%2C+22%7D 22, 25, 34, 35, 41, 41, 46, 46, 46, 47, 49, 54, 54, 59, 60 q0 = 22 q4 = 60 those are just the end values, the q1, q2, q3 are the meat of it

OpenStudy (amistre64):

if we get a decimal, we are lucky; just rnd up 3.75 -> 4 7.5 -> 8 11.25 -> 12 4 8 12 22, 25, 34, 35, 41, 41, 46, 46, 46, 47, 49, 54, 54, 59, 60 these are the box and whisker parts

OpenStudy (amistre64):

got no idea where "your question" is ... but if you ta me in it ill find it better